A sample of a refrigeration gas in a volume of 455 mL, at a pressure of 1.51 atm, and at a temperature of 25.0C was compressed into a volume of 220 mL with a pressure of 2.00 atm. To what temperature (in C) did it have to change?


1
Expert's answer
2021-03-08T07:07:06-0500

pV/T= p1V1/T1.

T=pVT1/p1V1 = 2 x 220 x 298/1.51 x 445 = 177,394151 K. = -95,605849 °©